FLUORESCENCE DEPLETION SPECTROSCOPY AND PHOTODYNAMICS OF THE Ã^{2}E ELECTRONIC STATE OF THE CADMIUM MONOMETHYL RADICAL

\underline{MICHAEL B. PUSHKARSKY} and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

The fluorescence depletion spectroscopy (FDS) technique has been applied to the Ã^{2}E electronic state of the CdCH_{3} radical to investigate “dark” vibronic levels. The observed FDS spectrum complements the sparse laser induced fluorescence (LIF) spectrum. Both the upper and lower spin components of the vibrationless level as well as excitations involving the symmetric ν_{2} and ν_{3} modes, and the antisymmetric ν_{6} mode in combination with both symmetric modes, were observed. A combined Jahn-Teller and spin-orbit coupling model was used to explain the vibronic structure in the ^{2}E electronic state as well as strong variations in the Coriolis coupling constants obtained from the resolved rotational K-structure. Lifetimes for fluorescing levels have been determined from their temporal decay and those for some of the dark levels via lifetime broadening measurements. The selectivity of the radiationless decay channel and it’s possible attribution to an isomerization process will be discussed.